OpenStudy (darkigloo):

Let y=f(x) be a particular solution to the differential equation dy/dx=xy^3 with f(1)=2. Write an equation for the tangent line to the graph of y=f(x) at x=1.

OpenStudy (darkigloo):

should this be my first step? int (y^-3)dy = int x dx

OpenStudy (loser66):

I think so

OpenStudy (darkigloo):

and then would i have to find c then plug in c to find y?

OpenStudy (loser66):

after take integral both sides , what do you have?

OpenStudy (darkigloo):

\[ \frac{ -1}{ 2y^2 }=\frac{ x^2 }{ 2} + C\]

OpenStudy (loser66):

solve for y,

OpenStudy (darkigloo):

\[(\frac{ -2 }{ x^2}+ \frac{ 1 }{ C } ) / 2 = y ^{2}\]

OpenStudy (loser66):

y =\(\pm \sqrt{thewholething}\) replace x =1 y =2 to solve for C

OpenStudy (darkigloo):

c=1/6?

OpenStudy (loser66):

actually, from the beginning, you have y =f(1) =2 mean point (1,2) is on the graph and you have to find out the tangent line at this point, right? One more thing, from the original problem , they give you dy /dx = xy^3 which is the slope of the tangent line of the curve defined by xy^3, so, at x =1, y=2, dy/dx =m = 8 you have the slope, and the point, just write out the tangent line by y = mx + b

OpenStudy (loser66):

solving ODE in the case they ask another question for equation of the curve. If they just ask the tangent line, do without solving the ODE , you still have the result.

OpenStudy (darkigloo):

ohh ok. so the answer is f(x)=2+8(x-1)

OpenStudy (loser66):

yup, should make it clear by \[\Large y_{tangent line} = 8x -6\]

OpenStudy (darkigloo):

ok thank you!